What value of x > -1 maximizes the integral ∫_(-1)^x▒〖t^2 (3-t) 〗 dt?


1
Expert's answer
2021-04-15T06:56:40-0400
"g(x)=\displaystyle\int_{-1}^xt^2(3-t)dt"


By the Fundamental Theorem of Calculus part I


"g'(x)=x^2 (3-x)"


Find the critical number(s)


"g'(x)=0, x>-1""x^2 (3-x)=0""x_1=0, x_2=3"


Critical numbers: "0, 3."

If "-1<x<0, g'(x)>0 , g(x)"  increases.

If "0<x<3, g'(x)>0 , g(x)"  increases.


If "x>3, g'(x)<0 , g(x)" decreases.

The function "g(x)" has a local maximum at "x=3."

Since the function "g(x)" has the only extremum for "x>-1," then the function "g(x)" has the absolute maximum at "x=3."


The integral "\displaystyle\int_{-1}^xt^2(3-t)dt" has the absolute maximum at "x=3."